(15 pt) A real number is represented by the IEEE floating format in 4-bit exponent and 6-bit fraction (ignore the sign bit). 9. What is the floating point representation of the number 53/11 in the format ? a. Exponent Fraction b. What are values of the following floating-point numbers in binary scientific notations (e.g. 1.10*29 ? 0000 000001 0101 010101
Answer #1

a. 53/11 = 4.81818181 = 100.11010001011101000112

Normalizing the result, we get

= 1.001101000101110100011 x 22

Exponent (biased) = 2 + 7 = 9

Fraction = 001101

b)

0000 000001 = 1.0 * 2-10

0101 010101 = 1.01010101 * 2-2